OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 2 ?

OpenStudy (mertsj):

log a - log b = log( a/b)

OpenStudy (mertsj):

Use that property first and then put the equation into its exponential form.

OpenStudy (owlcoffee):

so we have: \[\log_{2} x - \log_{2} 4=2\] Applying the property mertsj said, we'll end up with: \[\log_{2} \frac{ x }{ 4 }=2\] Now, we have to get rid of that logaritm so we'll transform the 2 on the right side of the equality to a logaritm that we know is true: \[\log_{2} 4= 2 \] so let's replace it: \[\log_{2} \frac{ x }{ 4 }=\log_{2} 4\] So now that we have Log2 on both sides, we can work like a very basic algebraic equation: \[\frac{ x }{ 4 }=4\] I encourage you to solve for x here and on :)

OpenStudy (anonymous):

x=16?

OpenStudy (owlcoffee):

Good, now to verify the value, replace 16 on the x of the orignal equation.